Omni Telecom is trying to decide whether to increase its cash dividend immediately or use the funds to increase its future growth rate. P0 = D1 Ke − g P0 = Price of the stock today D1 = Dividend at the end of the first year D1 = D0 × (1 + g) D0 = Dividend today Ke = Required rate of return g = Constant growth rate in dividends D0 is currently $2.00, Ke is 10 percent, and g is 4 percent. Under Plan A, D0 would be immediately increased to $2.50 and Ke and g will remain unchanged. Under Plan B, D0 will remain at $2.00 but g will go up to 5 percent and Ke will remain unchanged. a. Compute P0 (price of the stock today) under Plan A. Note D1 will be equal to D0 × (1 + g) or $2.50 (1.04). Ke will equal 10 percent, and g will equal 4 percent. (Round your intermediate calculations and final answer to 2 decimal places.) b. Compute P0 (price of the stock today) under Plan B. Note D1 will be equal to D0 × (1 + g) or $2.00 (1.05). Ke will be equal to 10 percent, and g will be equal to 5 percent. (Round your intermediate calculations and final answer to 2 decimal places.) c. Which plan will produce the higher value? Plan B Plan A
Answer #1

Solution:

Solution 1) Calculation of the value of stock under Plan A:

Required Rate of Return = Ke = 10%

Current Dividend = D0 = $2.50

Growth rate = g = 4%

Therefore, Expected Dividend = D1 = D0 (1 + g) = 2.50 (1.04) = $2.60

Value of Stock for Plan A = P0 = D1 / (Ke - g) = 2.60 / (0.10 – 0.04) = $43.33

Therefore, value of the stock under Plan A is $43.33.

       

Solution 2) Calculation of the value of stock under Plan B:

Required Rate of Return = Ke = 10%

Current Dividend = D0 = $2.00

Growth rate = g = 5%

Therefore, Expected Dividend = D1 = D0 (1 + g) = 2.00 (1.05) = $2.10

Value of Stock for Plan B = P0 = D1 / (Ke - g) = 2.10 / (0.10 – 0.05) = $42.00

Therefore, value of the stock under Plan B is $42.00.

Solution c) The company should go with Plan A as it is producing higher value for the stock i.e. $43.33